# On completeness of Foliated structures, and Null Killing fields

### v1

@@ 1. Introduction
Let $(M,\nabla)$ be a (connected) compact smooth manifold equipped with a smooth affine connection on the tangent bundle $TM$. A natural question to ask is: under which assumption on the couple $(M,\nabla)$ the manifold is geodesically complete (i.e. every geodesic is defined on $\mathbb{R}$)? Historically this started for $M$ having a smooth Riemannian metric $g$ and $\nabla$ being the associated Levi-Civita connection. Such a couple is always geodesically complete. However, for a non-Riemannian metric, the analogous statement is not true in general. Thus, the study is more rich in this direction, especially in Lorentzian signature, by providing sufficient conditions (curvature conditions, existence of Killing (or parallel) fields,..) for a compact Lorentzian manifold to be geodesically complete [6, 21, 16, 17, 18].
@@
In this paper we introduce a foliated version of this completeness problem. Namely, we consider a triple $(M,\mathfrak{F},\nabla_{\mathfrak{F}})$, where $M$ is a compact manifold, $\mathfrak{F}$ a foliation of $M$, and $\nabla_{\mathfrak{F}}$ a smooth affine connection defined on $T\mathfrak{F}$. The completeness problem above can be seen as the codimension 00 case in this setting. We formulate the problem as follows
@@
Of course, when the connection on a leaf of the foliation is the Levi-Civita connection of some Riemannian metric on the leaf, the geodesics contained in this leaf are complete. We refer to this as ‘a Riemannian situation’: this is an elementary situation from the point of view of the completeness problem. So we are interested in non-Riemannian situations. We consider in the following question a natural and interesting non-Riemannian situation:
@@
Let $(M,\mathfrak{F},\nabla_{\mathfrak{F}})$ be a triple as above. Assume that we have a degenerate Riemannian metric defined on the tangent bundle $T\mathfrak{F}$, such that on each leaf, the induced degenerate Riemannian metric is parallel with respect to the existing connection. Assume further that the kernel of the metric has dimension $1$, and denote by $\mathcal{N}$ the $1$-dimensional foliation it defines. Observe that the leaves of $\mathcal{N}$ are totally geodesic (indeed, the degenerate Riemannian metric on each leaf is parallel, so its kernel is invariant by parallel transport). Is it true that the leaves are complete if and only if the geodesics tangent to $\mathcal{N}$ are complete?
We do not give an answer in the full generality of Question 1.2, but this question contains many natural situations, which we study in this paper. One of them is that of a totally geodesic lightlike (i.e. the metric induced on each leaf is Riemannian degenerate) foliation $\mathfrak{F}$ of a compact Lorentzian manifold. In this case, the connection on the tangent bundle $T\mathfrak{F}$ is the restriction to $T\mathfrak{F}$ of the Levi-Civita connection of the Lorentzian metric. Another natural and important situation is when $\nabla$ is a flat connection. In this case, the leaves have some specific affine structure. The latter case actually belongs to a more general situation arising from Problem 1.1: that of a foliation with a $(G,X)$-structure on the leaves. We introduce the notion of a $(G,X)$-foliation in Definition 2.3, and we focus on certain affine structures with interesting application in the study of completeness in Lorentzian (or semi-Riemannian) geometry. In this case, the geodesic completeness is equivalent to the completeness of the structure when the model is complete. We refer to Subsection 1.3 of the introduction for the treatment of the problem in this setting.
@@
    A foliated compact manifold where the leaves have an affine structure in the sense of $(G,X)$-structures (see Definition 2.3). Observe that here, there is no assumption on existence of an ambient connection. The most important case of Item (B) we consider here is the case of the ‘affine lightlike geometry’ (see Subsection 1.3). The latter solves Question 1.2 when $\nabla$ is a flat connection.
@@
Theorem 1.3 contains the class of (compact) Brinkmann spacetimes. A Brinkmann spacetime is a Lorentzian manifold with a parallel null vector field $V$, so the codimension $1$ foliation $\mathfrak{F}$ tangent to $V^{\perp}$ is lightlike and totally geodesic. It is proved in [18] that a compact Brinkmann spacetime is complete. So this contains in particular a leafwise completeness result (of geodesics tangent to $\mathfrak{F}$). Theorem 1.3 can be seen as a generalization of the leafwise completeness to the situation where only a foliation is given, and this contains for instance the wider class of (compact) Kundt spacetimes (see [5, Proposition 2.1]).
As an application we get the following consequences
@@
In [17, Section 4.3, Example 3], the authors ask whether a compact Lorentzian manifold with a parallel null line field $l$ has a foliation (tangent to $l^{\perp}$) with complete leaves. The corollary above gives a necessary and sufficient condition for this to be true. Moreover, we give an example in which this condition does not hold, leading to incompleteness along the leaves (Example 3.5).
@@
Let $(M,V)$ be a compact Lorentzian manifold with a null Killing field. Assume that the $V^{\perp}$-distribution is integrable (i.e. involutive). Then the foliation tangent to $V^{\perp}$ is geodesically complete.
@@ 1.2. On the geodesic completeness in presence of a null Killing field
For a compact Lorentzian manifold with a null Killing field, one could in fact consider the question of global completeness, i.e. of the whole manifold $M$. This question is not treated in the literature. When the metric admits a timelike Killing field $V$, it was observed in [21] that the intersection of the levels of the metric with the levels of $g(\cdot,V)$ are compact, which leads to completeness. However, this is no longer the case when $V$ is lightlike or spacelike. In fact, in the spacelike case, taking a product of a non-complete Lorentzian manifold with the circle (with a Riemannian metric) shows that existence of a spacelike Killing field does not imply completeness. But there is no known incomplete example in the lightlike case.
@@ 1.2.1. An incomplete example
We construct in this paper two examples of incomplete compact Lorentzian manifolds with a null Killing field. The first one is locally homogeneous, given by the quotient of a left invariant metric on $G:={\mathsf{Sol}}\times\mathbb{R}$. We state
@@
There exist an incomplete left invariant Lorentzian metric on $G:={\mathsf{Sol}}\times\mathbb{R}$ with a central null Killing field $V$. Taking a quotient $M:=\Gamma\backslash G$, where $\Gamma$ is a cocompact lattice of $G$, defines an incomplete compact Lorentz manifold on which the flow of $V$ acts isometrically.
@@ 1.2.2. The case of dimension $3$
It is an easy fact that a compact Lorentzian surface with a null Killing field is complete (it is even flat). Our second incomplete example is more elaborate and is constructed in dimension $3$. Let $(M,V)$ be a compact Lorentzian manifold of dimension $3$, with $V$ a lightlike Killing field. The first observation is that in this dimension, the geodesics orthogonal to $V$ are complete. For the global completeness question, it turns out that the situation is different, depending on the dynamics of $V$: when $V$ has complicated dynamics (non-equicontinuous), the manifold is complete; in the opposite situation (equicontinuous case), we construct an incomplete example (see paragraphs below).
@@
In this case, the closure of the flow of $V$ in the isometry group is non-compact. The global behaviour of $V$ is well understood [25]: Zeghib classifies all such flows on Lorentzian $3$-manifolds. In particular, he proves that $V^{\perp}$ is integrable and geodesic. And it follows from the classification that $M$ is complete [25, Remark 16.1]. Completeness for geodesics transversal to $V^{\perp}$ seems non-obvious in general, and the question is still open in higher dimension.
@@
The case where the flow of $V$ is equicontinuous is still completely unknown. In this case, the closure of the flow of $V$ in the isometry group is an abelian compact subgroup, hence a torus of some dimension $k$. We show that (Proposition 7.3) if $k\geq 3$, the manifold is just a flat $3$-torus. When $k=2$, we show completeness in some special situations, namely when the $\mathbb{T}^{2}$-action is free and the fibers of the $\mathbb{T}^{2}$-principal bundle have the same signature. Finally, we focus on the case where $k=1$. In this case, the action of $V$ is periodic, and when it is free, $M$ is an $\mathbb{S}^{1}$-bundle over a torus. The study of this situation raises interesting general questions on the relationship between the geodesic completeness and the dynamics of certain vector fields defined on the torus (we hope to come back to it in a future work).
@@
There exist an incomplete compact $3$-dimensional Lorentzian manifold with a null Killing field.
The next subsection is devoted to the second part (B).
@@ 1.3. $(G,X)$-foliated completeness
Completeness of compact $(G,X)$-manifolds is a very rich and active topic of old and current research. It has been developed in many directions, especially in the case of affine structures, for instance [6, 9, 8, 3]. Completeness of compact $(G,X)$-manifolds is in general far from being easy. Here we consider a compact manifold $M$ with a $(G,X)$-foliation $\mathfrak{F}$, i.e. a geometric structure supported on a foliation $\mathfrak{F}$ of $M$. We call the couple $(M,\mathfrak{F})$ a $(G,X)$-foliated manifold. Roughly speaking, the leaves have a $(G,X)$-structure that varies continuously from a leaf to another. Those manifolds arise naturally in many interesting geometric settings. For example, the so-called pp-waves (Example 2.4), and more generally, weakly pp-waves (Example 3.1), are codimension $1$ foliated Lorentzian manifolds with a rigid structure on the leaves. Namely, the leaves of a pp-wave (resp. of a weakly pp-wave) have a special affine structure called “the affine unimodular lightlike geometry” (resp. “the affine lightlike geometry”), introduced in subsections 1.3.1 and 1.3.2 below. Note that $(G,X)$-foliations are briefly mentioned in Thurston’s book [22, p.160], but not precisely defined. Since they arise naturally, it is interesting to consider their completeness. A more precise formulation of Item (B) is the following
@@
Consider a $(G,X)$-structure for which any compact $(G,X)$-manifold is complete, we ask the following question: for a compact $(G,X)$-foliated manifold $M$, are the leaves $(G,X)$-complete? We call this the $(G,X)$-foliated completeness problem.
To the best of our knowledge, the $(G,X)$-foliated completeness problem has not been considered before. Besides the fact that it is itself interesting and natural, there are some particular geometric structures for which the full completeness reduces essentially to a $(G,X)$-foliated completeness problem.
@@ 1.3.1. Affine unimodular lightlike geometry
An $(\sf{Aff}(\mathbb{R}^{n}),\mathbb{R}^{n})$-structure on a manifold $M$ is called an affine structure. For a complete survey on affine structures one can see [10, Ch. 6-10]. We will be interested in a rigid version of the full affine geometry, which we call the affine unimodular lightlike geometry. It was first introduced in [14, page 3]. The structure group is a subgroup of the affine group, given by
@@
where $\mathsf{O}(n)$ is the orthogonal group of the Euclidean scalar product on $\mathbb{R}^{n}$. The $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry is the maximal affine geometry with two $\mathsf{L}_{\sf{u}}(n,0)$-invariant objects: the vector field $V:=\partial_{x_{1}}$ and a degenerate Riemannian metric $h:=dx_{2}^{2}+dx_{3}^{2}+..+dx_{n+1}^{2}$, with kernel $V$. The notation $(n,0)$ refers to the signature of this degenerate Riemannian metric on $\mathbb{R}^{n+1}$, and is useful to distinguish between different signatures in the next sections. However, when the second entry is 00, we simply denote ${\sf{L}}_{\sf{u}}(n)$ instead of $\mathsf{L}_{\sf{u}}(n,0)$.
It turns out that the (flat) connection preserved by ${\sf{L}}_{\sf{u}}(n)$ belongs to what is called in [8] distal connections. It is proved in the same paper that such a connection is complete when the manifold is compact. For the foliated problem (B), we show
@@
This gives for this geometry an affirmative answer to $(G,X)$-foliated completeness problem.
One could also consider a “dual” structure to $({\sf{L}}_{\sf{u}}(n),\mathbb{R}^{n+1})$, where the structure group is defined by
@@
The linear part of the structure group ${\sf{L}}_{\sf{u}}^{*}(n)$ is the dual of the linear part of ${\sf{L}}_{\sf{u}}(n)$, hence the name of dual geometry. It has an invariant parallel $1$-form (instead of an invariant parallel vector field) and an invariant foliation (instead of only a distribution). For this dual geometry, we establish an analogue statement to Theorem 1.9.

### v2

# and Null Killing fields

@@ Abstract.
Mathematics Subject Classification: 53C12, 53C50, 53C22.
Malek Hanounah
Institut für Mathematik und Informatik
Greifswald University, Germany
Email: malek.hanounah@uni-greifswald.de
Lilia Mehidi
Departamento de Geometria y Topologia
Facultad de Ciencias, Universidad de Granada, Spain
Email: lilia.mehidi@ugr.es
Statements and Declarations
Competing interests:
The authors have no competing interests to declare that are relevant to the content of this article.
Data availibility Data sharing is not applicable to this article as no new data were created or analyzed in this study.
@@ 1. Introduction
Let $(M,\nabla)$ be a (connected) compact smooth manifold equipped with a smooth affine connection on the tangent bundle $TM$. A natural question is: under what conditions on the pair $(M,\nabla)$ is the manifold geodesically complete, meaning every geodesic is defined on $\mathbb{R}$? When $\nabla$ is the Levi-Civita connection of a smooth Riemannian metric $g$ on $M$, such a couple is geodesically complete. However, for non-Riemannian metrics, for instance in Lorentzian signature, this statement does not hold in general, making the study more rich. This leads to exploring sufficient conditions, such as curvature constraints, or the existence of Killing (or parallel) vector fields, for a compact Lorentzian manifold to be geodesically complete [6, 32, 21, 23, 26].
@@
In various settings, one may naturally ask if some classical results extend to the context of foliations, i.e. whether they have a ‘foliated’ counterpart. A well-known example is the uniformization problem of $2$-dimensional foliations [12] (or more generally, laminations), which asks whether a Riemannian metric $g$ defined on the tangent bundle of a codimension $1$ foliation (or lamination) of a compact manifold of dimension $3$, can be made conformal to a metric $h$ with constant curvature along the leaves. Another example asks whether leafwise harmonic functions for foliations on compact manifolds have the same behavior as harmonic functions on compact manifolds [7]. In this paper, we introduce foliated versions of completeness problems. Namely, we consider a triple $(M,\mathfrak{F},\nabla_{\mathfrak{F}})$, where $M$ is a compact manifold, $\mathfrak{F}$ a foliation of $M$, and $\nabla_{\mathfrak{F}}$ a smooth affine connection defined on $T\mathfrak{F}$. The completeness problem above can be seen as the codimension 00 case in this setting. We formulate the problem as follows
@@
Of course, when the connection on the bundle $T\mathfrak{F}$ is induced by the Levi-Civita connection of a Riemannian metric on the manifold, the leaves are complete. We refer to this as a ‘Riemannian situation’: this is an elementary situation from the point of view of the completeness problem. Therefore, we are interested in non-Riemannian situations. In the following question, we consider a natural and interesting non-Riemannian situation:
@@
Let $(M,\mathfrak{F},\nabla_{\mathfrak{F}})$ be a triple as defined above. Assume that we have a degenerate Riemannian metric on the tangent bundle $T\mathfrak{F}$, such that on each leaf, the induced degenerate Riemannian metric is parallel with respect to the given connection. Assume further that the radical (kernel) of the metric has dimension $1$, and denote by $\mathcal{N}$ the $1$-dimensional foliation it defines. Observe that the leaves of $\mathcal{N}$ are totally geodesic (indeed, the degenerate Riemannian metric on each leaf is parallel, so its radical is invariant by parallel transport). Is it true that the leaves are complete if and only if the geodesics tangent to $\mathcal{N}$ are complete?
We do not give an answer to Question 1.2 in full generality, but this question contains many natural situations, which we study in this paper. One of them is that of a totally geodesic lightlike (i.e. the metric induced on each leaf is Riemannian degenerate) foliation $\mathfrak{F}$ of a compact Lorentzian manifold. In this case, the connection on the tangent bundle $T\mathfrak{F}$ is the restriction to $T\mathfrak{F}$ of the Levi-Civita connection of the Lorentzian metric. Another natural and important situation is when $\nabla$ is a flat connection. In this case, the leaves have some specific affine structure. The latter case actually belongs to a more general situation arising from Problem 1.1: that of a foliation with a $(G,X)$-structure on the leaves. We introduce the notion of a $(G,X)$-foliation in Definition 2.3, and we focus on certain affine structures with interesting application in the study of completeness in Lorentzian (or semi-Riemannian) geometry. In this case, the geodesic completeness is equivalent to the completeness of the $(G,X)$-structure when the model is complete. We refer to Subsection 1.3 of the introduction for the treatment of the problem in this setting.
@@
    A foliated compact manifold where the leaves have an affine structure in the sense of $(G,X)$-structures (see Definition 2.3). Observe that here, there is no assumption on existence of an ambient connection. The assumptions in Question 1.2, combined with the situation in Item (B), give rise to the so-called ‘affine lightlike geometry’ (see Subsection 1.3) on the leaves. We characterize the completeness in this case, providing an answer to Question 1.2 when $\nabla$ is a flat connection. Finally, we present other natural affine structures where geodesic completeness may fail.
@@
Theorem 1.3 includes the class of (compact) Brinkmann spacetimes. A Brinkmann spacetime is a Lorentzian manifold with a parallel null vector field $V$, so the codimension $1$ foliation $\mathfrak{F}$ tangent to $V^{\perp}$ is lightlike and totally geodesic. It is proved in [26] that a compact Brinkmann spacetime is complete. In particular, the completeness of geodesics tangent to $\mathfrak{F}$ is a leafwise completeness result. Theorem 1.3 can be viewed as a generalization of the leafwise completeness to the situation where only a totally geodesic foliation is given. This includes, for instance, the broader class of compact (locally) Kundt spacetimes (see [27] for a recent introductory survey).
As an application, we get the following consequences:
@@
In [23, Section 4.3, Example 3], the authors ask whether a compact Lorentzian manifold with a parallel null line field $l$ has a foliation (tangent to $l^{\perp}$) with complete leaves. The corollary above gives a necessary and sufficient condition for this to be true. In addition, we give an example where this condition is not satisfied, resulting in incomplete leaves. This example is constructed as the product of a circle with a Clifton-Pohl torus (Example 3.4).
@@
Let $(M,V)$ be a compact Lorentzian manifold with a null Killing field. Assume that the distribution $V^{\perp}$ is integrable. Then, the foliation tangent to $V^{\perp}$ is totally geodesic, and a geodesic tangent to it is complete.
@@ 1.2. On the geodesic completeness in presence of a null Killing field
Given a compact Lorentzian manifold with a null Killing field, one could extend the question to global completeness, that is, the completeness of the entire manifold $M$. One motivation for this is the completeness result obtained in [26] when $V$ is additionally parallel.
When the metric admits a timelike Killing field $V$, it was observed in [32] that the intersection of the levels of the metric with the levels of $g(\,\cdot,V)$ are compact, which leads to completeness. However, this is no longer the case when $V$ is spacelike or lightlike. In fact, in the spacelike case, taking a product of a non-complete Lorentzian manifold with the circle (endowed with a Riemannian metric) shows that existence of a spacelike Killing field does not imply completeness. But there is no known incomplete example in the lightlike case.
@@ 1.2.1. An incomplete example
We construct in this paper two examples of incomplete compact Lorentzian manifolds with a null Killing field. The first one is locally homogeneous. We state
@@
There exists an incomplete left invariant Lorentzian metric on $G:={\mathsf{Sol}}\times\mathbb{R}$ with a central null Killing field $V$. Taking a quotient $M:=\Gamma\backslash G$, where $\Gamma$ is a cocompact lattice of $G$, defines an incomplete compact Lorentzian manifold on which the flow of $V$ acts isometrically.
@@ 1.2.2. The case of dimension $3$
It is a simple fact that a compact Lorentzian surface with a null Killing field is complete (in fact, it is even flat). Our second incomplete example is more elaborate and is constructed in dimension $3$. Let $(M,V)$ be a compact Lorentzian manifold of dimension $3$ with $V$ a null Killing field. The first observation is that, in this dimension, the distribution $V^{\perp}$ is integrable and totally geodesic, and the geodesics tangent to $V^{\perp}$ are complete by Corollary 1.5. However, for the question of global completeness, the situation is different depending on the dynamics of $V$: when $V$ has complicated (non-equicontinuous) dynamics, the manifold is complete; in the opposite (equicontinuous) case, we construct an incomplete example (see details below).
@@
This means that the closure of the flow of $V$ in the isometry group is non-compact. The global behavior of $V$ is well understood [36]: Zeghib classifies all such flows on Lorentzian $3$-manifolds, and it follows from the classification that $M$ is complete [36, Remark 16.1]. The case where $V$ is non-equicontinuous leads to a non-compact isometry group. In this more general case, namely, when the isometry group of a compact Lorentzian $3$-manifold is non-compact, the manifold is complete by Frances’s classification results [8]. The question remains open in higher dimensions, and completeness of geodesics transversal to $V^{\perp}$ seems non-trivial in general.
@@
The case where the flow of $V$ is equicontinuous remains widely open. In this case, the closure of the flow of $V$ in the isometry group is an abelian compact subgroup, hence a torus of some dimension $k$. We show that if $k\geq 3$, the manifold is just a flat $3$-torus (Proposition 7.1). When $k=2$, we show completeness in some specific situations, namely, when the $\mathbb{T}^{2}$-action is free and the metrics induced on the fibers of the $\mathbb{T}^{2}$-principal bundle have the same signature. Finally, we consider the case where $k=1$. Here, the action of $V$ is periodic, and when it is free, $M$ is an $\mathbb{S}^{1}$-bundle over a torus. Even if the dynamics of $V$ is trivial, the reduction of the geodesic equation to dimension $2$, due to the existence of a Noether first integral, leads to a complicated situation. The study of this situation raises interesting questions on the relationship between the geodesic completeness and the dynamics of certain vector fields defined on the torus. We prove the following
@@
There exists an incomplete compact $3$-dimensional Lorentzian manifold with a periodic null Killing field.
The next subsection is devoted to Part (B).
@@ 1.3. $(G,X)$-foliated completeness
Completeness of compact $(G,X)$-manifolds is a very rich and active topic of old and current research. It has developed in many directions, especially in the case of affine structures (see, for instance, [6, 10, 9]). Completeness of compact $(G,X)$-manifolds is in general far from being easy. Here, we consider a compact manifold $M$ with a foliation $\mathfrak{F}$ having a tangential $(G,X)$-structure, meaning that the $(G,X)$-structure is supported on the leaves of $\mathfrak{F}$ and varies continuously from one leaf to another. We call the pair $(M,\mathfrak{F})$ a $(G,X)$-foliated manifold. A precise definition is given in Definition 2.3.
Note that $(G,X)$-foliated manifolds arise naturally in many interesting geometric contexts. For example, the so-called pp-waves (Example 1.9), and more generally, weakly pp-waves (Example 1.12), are codimension $1$ foliated Lorentzian manifolds with a rigid structure on the leaves. For a pp-wave (resp. weakly pp-wave), this structure is a particular affine structure called the ‘affine unimodular lightlike geometry’ (resp. the ‘affine lightlike geometry’). Those structures are introduced in Subsections 1.3.1 and 1.3.2 below.
Note that $(G,X)$-foliations are briefly mentioned in Thurston’s book [33, p.160], but not precisely defined. However, there is some old literature dealing with tangential structures on foliations. For instance, [11, 19] study foliations with an affine tangential structure. While these works address different questions (in [11] the author assumes completeness of the leaves to prove topological results on the manifold, while [19] deals with the existence of such foliations on the $2$-torus and the $3$-sphere, independently of completeness), the present paper introduces a new perspective by focusing on the completeness problem. A more precise formulation of Item (B) is the following
@@
Consider a $(G,X)$-structure for which any compact $(G,X)$-manifold is complete. Is it true that for any compact $(G,X)$-foliated manifold $(M,\mathfrak{F})$, the leaves of $\mathfrak{F}$ are $(G,X)$-complete? We call this the $(G,X)$-foliated completeness problem.
To our knowledge, the $(G,X)$-foliated completeness problem has not been considered before. Besides the fact that it is itself interesting and natural, there are some particular geometric structures for which the full completeness reduces essentially to a $(G,X)$-foliated completeness problem.
@@ 1.3.1. Affine unimodular lightlike geometry
An $(\sf{Aff}(\mathbb{R}^{n+1}),\mathbb{R}^{n+1})$-structure on a manifold $M$ is called an affine structure. For a complete survey on affine structures, one can see [14, Ch. 6-10]. We will be interested in a subgeometry of the affine geometry, which we call the affine unimodular lightlike geometry. It was first introduced in [17]. The structure group is a subgroup of the affine group $\sf{Aff}(\mathbb{R}^{n+1})$, given by
@@
where $\mathsf{O}(n)$ is the orthogonal group of the Euclidean scalar product on $\mathbb{R}^{n}$. The $(\mathsf{L}_{\sf{u}}(n,0),\mathbb{R}^{n+1})$-geometry is the maximal affine geometry with two $\mathsf{L}_{\sf{u}}(n,0)$-invariant objects: a constant vector field $V:=\partial_{x_{1}}$ and a degenerate Riemannian metric $h:=dx_{2}^{2}+dx_{3}^{2}+\ldots+dx_{n+1}^{2}$ with radical $\mathbb{R}V$. The notation $(n,0)$ refers to the signature of this degenerate Riemannian metric on $\mathbb{R}^{n+1}$, and is useful to distinguish between different signatures in the next sections. When the second entry is 00, we simply write ${\sf{L}}_{\sf{u}}(n)$ instead of $\mathsf{L}_{\sf{u}}(n,0)$.
A pp-wave is a Lorentzian manifold $(M,V)$ with a parallel null vector field $V$, such that the (integrable) parallel distribution $V^{\perp}$ is tangent to a (totally geodesic) foliation $\mathfrak{F}$ with flat leaves (flat with respect to the induced connection). It admits “Brinkmann local coordinates” of the form
$2dudv+H(u,x)du^{2}+\sum_{i=1}^{n}(dx^{i})^{2}.$
A pp-wave of dimension $n+2$, with its foliation $\mathfrak{F}$, provides a natural example of a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. Indeed, it is well known that the data of a flat affine connection on a smooth manifold of dimension $n+1$ is equivalent to having an $(\sf{Aff}(\mathbb{R}^{n+1}),\mathbb{R}^{n+1})$-structure on it. So $\mathfrak{F}$ inherits a tangential affine structure that varies smoothly. Moreover, the leaves of $\mathfrak{F}$ admit a tangent parallel lightlike vector field $V$, and an induced parallel degenerate Riemannian metric with radical $\mathbb{R}V$. This gives the leaves a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry.
It turns out that the (flat) connection preserved by ${\sf{L}}_{\sf{u}}(n)$ belongs to what is referred to as a distal connection in [9]. It is shown in this paper that such a connection is complete when the manifold is compact. For the foliated version, we show:
@@
This gives an affirmative answer to the $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated completeness problem.
One could also consider a “dual” structure to $({\sf{L}}_{\sf{u}}(n),\mathbb{R}^{n+1})$, where the structure group is
@@
The linear part of ${\sf{L}}_{\sf{u}}^{*}(n)$ is the dual of the linear part of ${\sf{L}}_{\sf{u}}(n)$, hence the name of dual geometry. It has an invariant parallel $1$-form (instead of an invariant parallel vector field), and an invariant foliation (instead of only a distribution). For this dual geometry, we establish an analogous statement to Theorem 1.10.